# VICON: Vision In-Context Operator Networks for Multi-Physics Fluid Dynamics Prediction

### v4

@@ 1 Introduction
Machine learning has emerged as a powerful tool for solving Partial Differential Equations (PDEs). Traditional approaches primarily operate on discrete representations, with Convolutional Neural Networks (CNNs) excelling at processing regular grids [1], Graph Neural Networks (GNNs) handling unstructured meshes [2, 3, 4], and transformers capturing dependencies between sampling points within the same domains [5, 6, 7]. However, these discrete approaches face a fundamental limitation: they lack PDE context information (such as governing parameters) and consequently cannot generalize to new parameters or PDEs beyond their training set.
@@
This limitation led to the development of “operator learning,” where neural networks learn mappings from input functions to output functions, such as from initial/boundary conditions to PDE solutions. This capability was first demonstrated using shallow neural networks [8, 9], followed by specialized architectures including the widely adopted Deep Operator Network (DeepONet) [10], Fourier Neural Operator (FNO) [11, 12], and recent transformer-based operator learners [13, 14, 15]. While these methods excel at learning individual parametric PDEs, they cannot generalize across different PDE types, necessitating costly retraining for new equations or even different timestep sizes.
@@
Inspired by the success of large language models in multi-domain generation, researchers have explored “multi-physics PDE models” to address this single-operator limitation. Common approaches employ pre-training and fine-tuning strategies [16, 17, 18], though these require substantial additional data (typically hundreds of frames) and fine-tuning before deployment. This requirement severely limits their practical applications in scenarios like online control and data assimilation, where rapid adaptation with limited data sampling is essential.
@@
To eliminate the need for additional data collection and fine-tuning, several works have attempted to achieve few-shot or zero-shot generalization across multi-physics systems by incorporating explicit PDE information. For example, the PROSE approach [19, 20, 21, 22] processes the symbolic form of governing equations using an additional transformer branch. Similarly, PDEformer [23] converts PDEs into directed graphs as input for graph transformers, while others embed symbol-tokens directly into the model [24]. Despite these advances, such methods require explicit knowledge of the underlying PDEs, which is often unavailable when deploying models in new environments.
@@
To further lower barriers for applying multi-physics models in real-world applications, the In-Context Operator Network (ICON) [25] offers a fundamentally different approach: ICON implicitly encodes system dynamics through a few input-output function pairs, then extracts dynamics from these pairs in an in-context fashion. This approach enables few-shot generalization without retraining or explicit PDE knowledge, while the minimal required input-output pairs can be readily collected during deployment. A single ICON model has demonstrated success in handling both forward and inverse problems across various ODEs, PDEs, and mean-field control scenarios.
@@
However, ICON’s computational efficiency becomes a critical bottleneck when handling dense data in higher dimensions. Specifically, ICON treats each sampled spatial point as an individual token, leading to quadratic computational complexity with respect to the number of points. This makes it computationally prohibitive for practical 2D and 3D applications—to date, only one instance of a 2D case using sparse data points has been demonstrated [25].
@@
To address this limitation, we propose Vision In-Context Operator Networks (VICON), which leverage vision transformers to process 2D functions using an efficient patch-wise approach. Going beyond classic vision transformers that typically process patches from a single image, VICON extends the architecture to handle sequences of input-output function pairs, enabling “next function prediction” capabilities while preserving the benefits of in-context operator learning.
@@
  • Comprehensive evaluation across diverse fluid dynamics systems comprising 879K frames with varying timestep sizes, demonstrating substantial improvements over state-of-the-art models MPP [26] and DPOT [27] in both accuracy and efficiency. Our approach reduces the averaged last-step rollout error by 37.9% compared to DPOT and 44.7% compared to MPP, while requiring only 72.5% and 34.8% of their respective inference times.
@@ 2 Related Work
Operator learning [8, 9, 11, 10] addresses the challenge of approximating operators $G:U\to V$, where $U$ and $V$ are function spaces representing physical systems and differential equations, e.g., $G$ maps initial/boundary conditions or system parameters to corresponding solutions. Among the various operator learning approaches, DeepONets [10, 28] employ branch and trunk networks to independently process inputs and query points, while FNOs [11] leverage fast Fourier transforms to efficiently compute kernel integrations for PDE solutions in regular domains. These methods have been extended to incorporate equation information [29, 30], multiscale features [31, 32, 33], and adaptation to heterogeneous and irregular meshes [34, 35, 36, 37].
@@
Foundation models in natural language processing [38, 39] and computer vision [40] have demonstrated remarkable versatility across diverse tasks. Drawing inspiration from this paradigm, recent research has explored developing unified models for multiple PDEs in domains such as PDE discovery [41] and computational fluid dynamics [42]. Several approaches follow pre-training and fine-tuning strategies [16, 17], though these require additional resources for data acquisition and fine-tuning before deployment. In parallel, researchers have developed zero- or few-shot multi-physics models by incorporating additional information; for instance, the PROSE approach [19, 20, 21, 22, 43] directly encodes symbolic information of PDEs into the model. Other methods achieving zero- or few-shot generalization include using physics-informed tokens [24], representing PDE structures as graphs [23], or conditioning transformers on PDE descriptions [44]. However, all these techniques require prior knowledge of the PDEs as additional inputs. We select two state-of-the-art models, MPP [26] and DPOT [27], as our baselines.
@@
ICONs [25, 45, 46] learn operators by observing input-output function pairs, enabling few-shot generalization across various PDEs without requiring explicit PDE representations or fine-tuning. These networks have demonstrated success in handling forward and inverse problems for ODEs, PDEs, and mean-field control scenarios. However, ICONs face significant computational challenges when processing dense data, as they represent functions through scattered point tokens, resulting in quadratic computational complexity with respect to the number of sample points. This limitation has largely restricted their application to 1D problems, with only sparse sampling feasible in 2D cases [25]. Our work addresses these limitations by introducing a vision transformer architecture that efficiently handles dense 2D data while preserving the benefits of in-context operator learning.
@@ 3 Preliminaries
ICONs were introduced and developed in [25, 45, 46]. They adapt the in-context learning framework from large language models, aiming to train a model that can learn operators from prompted function examples.
@@
For training parallelization, ICON uses a special causal attention mask to perform autoregressive learning (i.e., the model only sees the leading $J$ pairs and ${\bm{c}}_{J+1}$ when predicting $\tilde{{\bm{q}}}_{J+1}$) and enables output of all predictions within a single forward pass [45]:
@@ 4.2 Vision In-Context Operator Networks
Inspired by the Vision Transformer (ViT) [47], we address these limitations by dividing the physical fields into patches, where each patch is flattened and projected as a token. This approach, which we call Vision In-Context Operator Networks (VICON), addresses the computational limitation of treating each spatial point as an individual token, which leads to quadratic computational complexity with respect to the number of points. This patch-wise approach has its forward process illustrated in fig.˜1(b) and detailed in the following.
@@
where $k=1,\ldots,N_{c}\text{ and }l=1,\ldots,N_{q}\text{ are the index of patches.}$
We inject two types of learnable positional encoding before feeding the embeddings into the transformer: (1) patch positional encodings to indicate relative patch positions inside the whole domain, denoted as $\mathbf{E}_{p}\in\mathbb{R}^{N_{p}\times d}$, where $N_{p}=\max\{N_{c},N_{q}\}$ and the encoding varies per patch (broadcast across points within each patch); (2) function positional encodings to indicate whether a patch belongs to a input function (using $\mathbf{E}_{c}\in\mathbb{R}^{I\times d}$) or output function (using $\mathbf{E}_{q}\in\mathbb{R}^{I\times d}$), as well as their indices in the sequence, where the encoding varies per condition/QoI in each pair (broadcast across patches and points of that condition/QoI):
@@
To support autoregressive prediction similar to eq.˜2, VICON employs an alternating-sized (in the case where $N_{c}\neq N_{q}$) block causal attention mask, as opposed to the conventional triangular causal attention mask as in mainstream generative large language models [38] and large vision models [48]. The mask is defined as follows:
@@
After obtaining the output tokens from eq.˜2, we extract the tokens corresponding to the input patch indices $\mathbf{c}_{i}^{1},\ldots,\mathbf{c}_{i}^{N_{c}}$ and denote them as $\tilde{{\bm{q}}}_{i}^{1},\ldots,\tilde{{\bm{q}}}_{i}^{N_{c}}$. These tokens are then projected back to the original physical space using a shared learnable linear function $g_{\psi}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{R_{x}\times R_{y}\times c}$:
@@ 4.3 Prompt Normalization
In this work, we simply set $\mathcal{N}_{c}=\mathcal{N}_{q}$, mainly because our maximum timestep stride $s_{\text{max}}=5$ is relatively small (i.e., the scale distribution does not change dramatically). Specifically, we computed the channel-wise mean $\mathbf{\mu}$ and standard deviation $\mathbf{\sigma}$ of $\{{\bm{c}}_{i}\}_{i=1}^{I}$ in each prompt, then used these values to normalize $\{\langle{\bm{c}}_{i},{\bm{q}}_{i}\rangle\}_{i=1}^{I}$ in that prompt. To avoid division by zero, we set a minimum threshold of $10^{-4}$ for the standard deviation.
@@ 4.4 Datasets and Data Augmentation
We evaluate on three fluid dynamics datasets representing different physical regimes: 1) PDEArena-Incomp [1] (incompressible Navier-Stokes equations), containing 2,496/608/608 trajectories (train/valid/test) with 56 timesteps each; 2) PDEBench-Comp-HighVis [49] (compressible Navier-Stokes with high viscosity), containing 40,000 trajectories of 21 timesteps each; and 3) PDEBench-Comp-LowVis [49] (compressible Navier-Stokes with numerically zero viscosity), containing 4,000 trajectories of 21 timesteps each. For PDEArena-Incomp, we follow the original data split in [1]. For PDEBench-Comp-HighVis and PDEBench-Comp-LowVis, we randomly split trajectories in 80%/10%/10% proportions for training/validation/testing, respectively. For all datasets during rollout, we set $I_{0}=10$, i.e., we predict trajectories up to the last timestep given the initial 10 frames.
The temporally homogeneous Markovian property (section˜4.1) enables natural data augmentation for training and flexible rollout strategy for inference (section˜4.5): by striding with larger timesteps to reduce the number of autoregressive steps in long-term prediction tasks: $\Delta t=s\Delta\tau_{i_{p}}\;|\;s=1,2,\dots,s_{\text{max}}$, where $\tau_{i_{p}}$ is the timestep size for recording in trajectory $i_{p}$. During training, for each trajectory in the training set, we first sample a stride size $s\sim\mathcal{U}\{1,\ldots,s_{\text{max}}\}$, then form $\langle\mathbf{u}_{t},\mathbf{u}_{t+s\Delta\tau}\rangle$ pairs for the corresponding operator. We illustrate this augmentation process in fig.˜1(a), where different strides are randomly sampled during dataloading.
@@ 4.5 Inference with Flexible Strategies
As mentioned in section˜4.4, VICON can make predictions with varying timestep strides. Given $I_{0}$ initial frames $\{\mathbf{u}_{i}\}_{i=1}^{I_{0}}$, we can form in-context example pairs $\{\langle\mathbf{u}_{i},\mathbf{u}_{{i+s}}\rangle\}_{i=1}^{I_{0}-s}$. For $j\geq 1$, we set the question condition (${\bm{c}}_{k}$ in eq.˜3) as ${\bm{u}}_{j}$ to predict ${\bm{u}}_{j+s}$, enabling $s$-step prediction.
@@
For single-step rollout, we simply follow the autoregressive procedure with $s=1$. Flexible-step rollout involves a more sophisticated approach: given a maximum prediction stride $s_{\text{max}}$, we first make sequential predictions with a gradually growing $s=1$ to $s_{\text{max}}$ using $\mathbf{u}_{I_{0}}$ as the question condition, obtaining $\left\{\mathbf{u}_{I_{0}+s}\right\}_{s=1}^{s_{\text{max}}}$. Using each frame in this sequence as a question condition, we then make consistent $s_{\text{max}}$-step predictions while preserving all intermediate frames. This flexible-step strategy follows the approach in [46].
This strategy is also applicable to imperfect measurement cases where partial input frames are missing, e.g., due to sensor device error. In this case, we can still form pairs from the remaining frames, with minor adjustments to the strategy generation algorithms, as shown in fig.˜1(c).

### v5

@@ Abstract
11footnotetext: Equal contribution.22footnotetext: Work partially done while at UCLA.
@@ 1 Introduction
Machine learning has emerged as a powerful tool for solving Partial Differential Equations (PDEs). Traditional approaches primarily operate on discrete representations, with Convolutional Neural Networks (CNNs) excelling at processing regular grids Gupta and Brandstetter (2022), Graph Neural Networks (GNNs) handling unstructured meshes Sanchez-Gonzalez et al. (2020); Lino et al. (2022); Cao et al. (2023), and transformers capturing dependencies between sampling points within the same domains Cao (2021); Dang et al. (2022); Jiang et al. (2023). However, these discrete approaches face a fundamental limitation: they lack PDE context information (such as governing parameters) and consequently cannot generalize to new parameters or PDEs beyond their training set.
@@
This limitation led to the development of “operator learning,” where neural networks learn mappings from input functions to output functions, such as from initial/boundary conditions to PDE solutions. This capability was first demonstrated using shallow neural networks Chen and Chen (1995b; a), followed by specialized architectures including the widely adopted Deep Operator Network (DeepONet) Lu et al. (2021), Fourier Neural Operator (FNO) Li et al. (2021); Kovachki et al. (2023), and recent transformer-based operator learners Li et al. (2022; 2023); Ovadia et al. (2024). While these methods excel at learning individual parametric PDEs, they cannot generalize across different PDE types, necessitating costly retraining for new equations or even different timestep sizes.
@@
Inspired by the success of large language models in multi-domain generation, researchers have explored “multi-physics PDE models” to address this single-operator limitation. Common approaches employ pre-training and fine-tuning strategies Chen et al. (2024); Herde et al. (2024); Zhang et al. (2024c), though these require substantial additional data (typically hundreds of frames) and fine-tuning before deployment. This requirement severely limits their practical applications in scenarios like online control and data assimilation, where rapid adaptation with limited data sampling is essential.
@@
To eliminate the need for additional data collection and fine-tuning, several works have attempted to achieve few-shot or zero-shot generalization across multi-physics systems by incorporating explicit PDE information. For example, the PROSE approach Liu et al. (2024c; b); Sun et al. (2024a); Jollie et al. (2024) processes the symbolic form of governing equations using an additional transformer branch. Similarly, PDEformer Ye et al. (2024) converts PDEs into directed graphs as input for graph transformers, while others embed symbol-tokens directly into the model Lorsung et al. (2024). Despite these advances, such methods require explicit knowledge of the underlying PDEs, which is often unavailable when deploying models in new environments.
@@
To further lower barriers for applying multi-physics models in real-world applications, the In-Context Operator Network (ICON) Yang et al. (2023a) offers a fundamentally different approach: ICON implicitly encodes system dynamics through a few input-output function pairs, then extracts dynamics from these pairs in an in-context fashion. This approach enables few-shot generalization without retraining or explicit PDE knowledge, while the minimal required input-output pairs can be readily collected during deployment. A single ICON model has demonstrated success in handling both forward and inverse problems across various ODEs, PDEs, and mean-field control scenarios.
@@
However, ICON’s computational efficiency becomes a critical bottleneck when handling dense data in higher dimensions. Specifically, ICON treats each sampled spatial point as an individual token, leading to quadratic computational complexity with respect to the number of points. This makes it computationally prohibitive for practical 2D and 3D applications—to date, only one instance of a 2D case using sparse data points has been demonstrated Yang et al. (2023a).
@@
To address this limitation, we propose Vision In-Context Operator Networks (VICON), which leverage vision transformers to process 2D functions using an efficient patch-wise approach. Going beyond classic vision transformers that typically process patches from a single image, VICON extends the architecture to handle sequences of input-output function pairs, enabling “next function prediction” capabilities while preserving the benefits of in-context operator learning. The current implementation focuses on passive, autonomous PDEs; extension to non-autonomous systems with external forcing is left for future work.
@@
  • Comprehensive evaluation across diverse fluid dynamics systems comprising 879K frames with varying timestep sizes, demonstrating substantial improvements over state-of-the-art models MPP McCabe et al. (2023) and DPOT Hao et al. (2024) in both accuracy and efficiency. Our approach reduces the averaged last-step rollout error by 37.9% compared to DPOT and 44.7% compared to MPP, while requiring only 72.5% and 34.8% of their respective inference times.
@@ 2 Related Work
Operator learning Chen and Chen (1995b; a); Li et al. (2021); Lu et al. (2021) addresses the challenge of approximating operators $G:U\to V$, where $U$ and $V$ are function spaces representing physical systems and differential equations, e.g., $G$ maps initial/boundary conditions or system parameters to corresponding solutions. Among the various operator learning approaches, DeepONets Lu et al. (2021; 2022) employ branch and trunk networks to independently process inputs and query points, while FNOs Li et al. (2021) leverage fast Fourier transforms to efficiently compute kernel integrations for PDE solutions in regular domains. These methods have been extended to incorporate equation information Wang et al. (2021); Li et al. (2024b), multiscale features Zhang et al. (2024a); Wen et al. (2022); Liu et al. (2024a), and adaptation to heterogeneous and irregular meshes Zhang et al. (2023; 2024b); Li et al. (2024a); Wu et al. (2024).
@@
Foundation models in natural language processing Brown et al. (2020); Touvron et al. (2023) and computer vision Ramesh et al. (2021) have demonstrated remarkable versatility across diverse tasks. Drawing inspiration from this paradigm, recent research has explored developing unified models for multiple PDEs in domains such as PDE discovery Schaeffer (2017) and computational fluid dynamics Wang et al. (2024a). Several approaches follow pre-training and fine-tuning strategies Chen et al. (2024); Herde et al. (2024), though these require additional resources for data acquisition and fine-tuning before deployment. In parallel, researchers have developed zero- or few-shot multi-physics models by incorporating additional information; for instance, the PROSE approach Liu et al. (2024c; b); Sun et al. (2024a); Jollie et al. (2024); Sun et al. (2024b) directly encodes symbolic information of PDEs into the model. Other methods achieving zero- or few-shot generalization include using physics-informed tokens Lorsung et al. (2024), representing PDE structures as graphs Ye et al. (2024), or conditioning transformers on PDE descriptions Zhou et al. (2025). However, all these techniques require prior knowledge of the PDEs as additional inputs. We select two state-of-the-art models, MPP McCabe et al. (2023) and DPOT Hao et al. (2024), as our baselines.
@@
ICONs Yang et al. (2023a; b); Yang and Osher (2024) learn operators by observing input-output function pairs, enabling few-shot generalization across various PDEs without requiring explicit PDE representations or fine-tuning. These networks have demonstrated success in handling forward and inverse problems for ODEs, PDEs, and mean-field control scenarios. However, ICONs face significant computational challenges when processing dense data, as they represent functions through scattered point tokens, resulting in quadratic computational complexity with respect to the number of sample points. This limitation has largely restricted their application to 1D problems, with only sparse sampling feasible in 2D cases Yang et al. (2023a). Our work addresses these limitations by introducing a vision transformer architecture that efficiently handles dense 2D data while preserving the benefits of in-context operator learning.
@@ 3 Preliminaries
ICONs were introduced and developed in Yang et al. (2023a; b); Yang and Osher (2024). They adapt the in-context learning framework from large language models, aiming to train a model that can learn operators from prompted function examples.
@@
For training parallelization, ICON uses a special causal attention mask to perform autoregressive learning (i.e., the model only sees the leading $J$ pairs and ${\bm{c}}_{J+1}$ when predicting $\tilde{{\bm{q}}}_{J+1}$) and enables output of all predictions within a single forward pass Yang et al. (2023b):
@@ 4.2 Vision In-Context Operator Networks
Inspired by the Vision Transformer (ViT) Dosovitskiy et al. (2021), we address these limitations by dividing the physical fields into patches, where each patch is flattened and projected as a token. This approach, which we call Vision In-Context Operator Networks (VICON), addresses the computational limitation of treating each spatial point as an individual token, which leads to quadratic computational complexity with respect to the number of points. This patch-wise approach has its forward process illustrated in Figure˜1(b) and detailed in the following.
@@
where $k=1,\ldots,N_{c}\text{ and }l=1,\ldots,N_{q}\text{ are the index of patches.}$ We use the same projection $f_{\phi}$ for both conditions and QoIs because in our setup, both represent the same physical fields at different timesteps; sharing the projection enforces consistency in latent space regardless of whether a frame serves as context or target.
We inject two types of learnable positional encoding before feeding the embeddings into the transformer: (1) patch positional encodings to indicate relative patch positions inside the whole domain, denoted as $\mathbf{E}_{p}\in\mathbb{R}^{N_{p}\times d}$, where $N_{p}=\max\{N_{c},N_{q}\}$ and the encoding varies per patch (broadcast across points within each patch); (2) function positional encodings to indicate whether a patch belongs to an input function (using $\mathbf{E}_{c}\in\mathbb{R}^{I\times d}$) or output function (using $\mathbf{E}_{q}\in\mathbb{R}^{I\times d}$), as well as their indices in the sequence, where the encoding varies per condition/QoI in each pair (broadcast across patches and points of that condition/QoI):
@@
To support autoregressive prediction similar to Equation˜2, VICON employs an alternating-sized (in the case where $N_{c}\neq N_{q}$) block causal attention mask, as opposed to the conventional triangular causal attention mask as in mainstream generative large language models Brown et al. (2020) and large vision models Bai et al. (2024). The mask is defined as follows:
@@
After obtaining the output tokens from Equation˜2, we extract the tokens corresponding to the input patch indices $\mathbf{c}_{i}^{1},\ldots,\mathbf{c}_{i}^{N_{c}}$ and denote them as $\tilde{{\bm{q}}}_{i}^{1},\ldots,\tilde{{\bm{q}}}_{i}^{N_{c}}$. These tokens are then projected back to the original physical space using a shared learnable linear function $g_{\psi}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{R_{x}\times R_{y}\times c}$:
@@ 4.3 Prompt Normalization
In this work, we simply set $\mathcal{N}_{c}=\mathcal{N}_{q}$, mainly because our maximum timestep stride $s_{\text{max}}=5$ is relatively small (i.e., the scale distribution does not change dramatically). Specifically, given input sequence $\{\langle{\bm{c}}_{i},{\bm{q}}_{i}\rangle\}_{i=1}^{I}$, we compute the channel-wise mean $\mathbf{\mu}=\texttt{mean}({\bm{c}}_{1},{\bm{c}}_{2},\dots,{\bm{c}}_{I})\in\mathbb{R}^{d}$ and standard deviation $\mathbf{\sigma}=\texttt{std}({\bm{c}}_{1},{\bm{c}}_{2},\dots,{\bm{c}}_{I})\in\mathbb{R}^{d}$ of all conditions $\{{\bm{c}}_{i}\}_{i=1}^{I}$, which are then used to normalize both the conditions and QoIs. Training MSE loss is computed in the normalized space. To avoid division by zero, we set a minimum threshold of $10^{-4}$ for the standard deviation.
@@ 4.4 Datasets and Data Augmentation
We evaluate on three fluid dynamics datasets representing different physical regimes: 1) PDEArena-Incomp Gupta and Brandstetter (2022) (incompressible Navier-Stokes equations), containing 2,496/608/608 trajectories (train/valid/test) with 56 timesteps each; 2) PDEBench-Comp-HighVis Takamoto et al. (2022) (compressible Navier-Stokes with high viscosity), containing 40,000 trajectories of 21 timesteps each; and 3) PDEBench-Comp-LowVis Takamoto et al. (2022) (compressible Navier-Stokes with numerically zero viscosity), containing 4,000 trajectories of 21 timesteps each. For PDEArena-Incomp, we follow the original data split in Gupta and Brandstetter (2022). For PDEBench-Comp-HighVis and PDEBench-Comp-LowVis, we randomly split trajectories in 80%/10%/10% proportions for training/validation/testing, respectively. For all datasets during rollout, we set $I_{0}=10$, i.e., we predict trajectories up to the last timestep given the initial 10 frames.
The temporally homogeneous Markovian property (Section˜4.1) enables natural data augmentation for training and flexible rollout strategy for inference (Section˜4.5): by striding with larger timesteps to reduce the number of autoregressive steps in long-term prediction tasks: $\Delta t=s\Delta\tau_{i_{p}}\;|\;s=1,2,\dots,s_{\text{max}}$, where $\tau_{i_{p}}$ is the timestep size for recording in trajectory $i_{p}$. During training, for each trajectory in the training set, we first sample a stride size $s\sim\mathcal{U}\{1,\ldots,s_{\text{max}}\}$, then form $\langle\mathbf{u}_{t},\mathbf{u}_{t+s\Delta\tau}\rangle$ pairs for the corresponding operator. We illustrate this augmentation process in Figure˜1(a), where different strides are randomly sampled during dataloading.
@@ 4.5 Inference with Flexible Strategies
As mentioned in Section˜4.4, VICON can make predictions with varying timestep strides. Given $I_{0}$ initial frames $\{\mathbf{u}_{i}\}_{i=1}^{I_{0}}$, we can form in-context example pairs $\{\langle\mathbf{u}_{i},\mathbf{u}_{{i+s}}\rangle\}_{i=1}^{I_{0}-s}$. For $j\geq 1$, we set the question condition (${\bm{c}}_{k}$ in Equation˜3) as ${\bm{u}}_{j}$ to predict ${\bm{u}}_{j+s}$, enabling $s$-step prediction.
@@
For single-step rollout, we simply follow the autoregressive procedure with $s=1$. Flexible-step rollout involves a more sophisticated approach: given a maximum prediction stride $s_{\text{max}}$, we first make sequential predictions with a gradually growing $s=1$ to $s_{\text{max}}$ using $\mathbf{u}_{I_{0}}$ as the question condition, obtaining $\left\{\mathbf{u}_{I_{0}+s}\right\}_{s=1}^{s_{\text{max}}}$. Using each frame in this sequence as a question condition, we then make consistent $s_{\text{max}}$-step predictions while preserving all intermediate frames. This flexible-step strategy follows the approach in Yang and Osher (2024).
This strategy is also applicable to imperfect measurement cases where partial input frames are missing, e.g., due to sensor device error. In this case, we can still form pairs from the remaining frames, with minor adjustments to the strategy generation algorithms, as shown in Figure˜1(c).